# On the reconstruction limits of complex networks

### v1

@@ III.1 Entropy
III.2 Network reconstructability
Let $M^{*}=(G^{*},X^{*})$ be the pair of random variables associated with the TDG model whose joint probability distribution is $P(G^{*})P(X^{*}|G^{*})$, and $M=(G,X)$ be some reconstruction model, which infers a graph $g$ with probability $P(G=g|X=x^{*})$ from a realisation $x^{*}$ of $X^{*}$. Also, let $I(G^{*};G|X=X^{*})\geq 0$ be the mutual information between the true and the reconstructed random graphs. The condition $X=X^{*}$ expresses the constraint that $X$ is set by realizations of $X^{*}$. This mutual information is defined as
$I(G^{*};G|X=X^{*})=H(G^{*}|X=X^{*})-H(G^{*}|G,X=X^{*})\,$ (10)
where $H(G^{*}|X=X^{*})=H(G^{*})$ by virtue of the fact that $G^{*}$ is conditionally independent of $X$ and
$H(G^{*}|G,X=X^{*})=-\mathbb{E}_{G^{*},X^{*},G}\quantity[\log P(G|\hat{G},X=X^{*})]\,$ (11)
is the entropy of the true graph as predicted by the reconstructed one, where $P(G^{*}|G,X=X^{*})=\frac{P(G^{*},G|X=X^{*})}{P(G|X=X^{*})}$ such that both $P(G^{*},G|X=X^{*})$ and $P(G|X=X^{*})$ are marginal distributions of $P(G^{*},X^{*},G|X=X^{*})$ [Eq. (2)]. The quantity $I(G^{*};G|X=X^{*})$ may be interpreted as measuring the similarity between the information contents of $G^{*}$ and $G$. The higher it is, the more similar $G^{*}$ and $G$ are and better is the reconstruction. Conversely, when $I(G^{*};G|X=X^{*})=0$, it is minimized and both graphs are independent from one another. Note that similar mutual information measures have been used as a performance measure in the context of community detection for comparing pairs of partitions Newman et al. (2020); Jerdee et al. (2024). Additionally, $I(G^{*};G|X=X^{*})$ is related to the probability of error, defined as
$p_{e}=P(G^{*}\neq G|X=X^{*})\,.$ (12)
This relationship can be shown through Fano’s inequality Cover and Thomas (2006):
$H(G^{*}|G,X=X^{*})\leq h(p_{e})+H(G^{*})p_{e}\,,$ (13)
where $h(p)\equiv-p\log p-(1-p)\log(1-p)$ is the binary entropy. Indeed, given that $h(p_{e})\leq 1$, Fano’s inequality can be modified such that
$p_{e}\geq 1-\frac{I(G^{*};G|X=X^{*})+1}{H(G^{*})}\,.$ (14)
This lower bound on the probability of error is minimized when $I(G^{*};G|X=X^{*})$ is maximized.
The data processing inequality Cover and Thomas (2006) stipulates that
$I(G^{*};G|X=X^{*})\leq I(G^{*};X^{*})\,.$ (15)
where the upper bound
@@ III.2 Network reconstructability
is the mutual information between the true graph $G^{*}$ and the data process $X^{*}$ (see the proof in Appendix D). Intuitively, $I(G^{*};X^{*})$ quantifies the amount of reconstructible information that both $X^{*}$ and $G^{*}$ share—i.e., the amount of information that $X^{*}$ contains about $G^{*}$ [see Fig. 1(b)]. The mutual information $I(G^{*};X^{*})$ also sets the maximum in reconstruction performance as measured by $I(G^{*};G|X=X^{*})$: It is the reconstruction limit.
The mutual information $I(G^{*};X^{*})$ is itself bounded between 00 and $H(G^{*})$ Cover and Thomas (2006). When $I(G^{*};X^{*})=0$, $X^{*}$ and $G^{*}$ are independent and thus the data $X^{*}$ contains no information about the graph $G^{*}$. In turn, it is impossible for any reconstruction model $M$ to extract information from the data, regardless of its specification. When $I(G^{*};X^{*})=H(G^{*})$, the data $X^{*}$ contains all the information about the graph $G^{*}$. In this case, it is in principle possible to perfectly reconstruct the graph without any error, assuming the model $M$ is optimal (we discuss the notion of optimality in Sec. III.3).
@@
The reconstructability has been described thoroughly in Ref. Murphy et al. (2024) and helped unveiling a special duality between our ability to predict the time evolution of a system and our ability to reconstruct the interactions between its constituents. As it is a normalized version of the mutual information upper bound $I(G^{*};X^{*})$, the reconstructability is bounded between 00 and $1$. When $U(G^{*}|X^{*})=0$, any attempt at reconstruction is futile, whereas it is theoretically possible to decode the all the information when $U(G^{*}|X^{*})=1$. As such, the reconstructability is a measure of the average proportion of information that can be extracted from the data about the graph. For instance, when it is equal to $\frac{1}{2}$, it precisely means that half of the graph information is, on average, contained in the data and that in turn half of it can possibly be reconstructed. We stress that $U(G|X)=\frac{1}{2}$ may not be directly interpreted as half of the graph’s edges being reconstructible. Rather, information may generally be distributed in a heterogeneous way over the graph’s structure, as a single bit of information may reconstruct more than one edge in the graph depending on how correlated they are.
Going back to the earlier example of a single edge of Sec. II.2, we can perform the complete calculation analytically. First, we can calculate the entropy of the prior,
@@
This leads to the reconstructability of the edge using Eq. (17), which is plotted in Fig. 3. As expected, the reconstructability typically increases as the number of observations $T$ increases, even reaching 1 in some cases, e.g. when $q\to 1$. Also, notice how the reconstructability is zero for every value of $T$ when the true positive probability $q$ is equal to false positive probability $r$. This shows that as true positives and false positives become indistinguishable, the edge becomes impossible to reconstruct.
III.3 Optimal reconstruction performance
The reconstruction limit corresponds to the maximum performance, as measured by $I(G^{*};G|X=X^{*})$, achievable by any algorithm. Hence, any model that is capable of reaching this limit, i.e., $I(G^{*};G|X=X^{*})=I(G;X)$, is said optimal in its reconstruction abilities. It may not be surprising that the TDG model is optimal according to this definition (proof available in Appendix D). As a result, the reconstructability $U(G^{*}|X^{*})$ may itself also be interpreted as a performance measure of the TDG model $M^{*}$.
The reconstructability is a meaningful performance metric of the TDG model that relates to other standard measures. One such example is the posterior loss—also known as the log loss and the cross-entropy loss in the machine learning community. This measure is defined as
@@ III.3 Optimal reconstruction performance
where $\bm{a}^{*}$ denotes the adjacency of the true graph $g^{*}$, such that $a_{ij}^{*}$ counts the number of edges connecting the nodes $i$ and $j$ (we use the convention that $a_{ii}^{*}$ is always a multiple of 2) and $\bm{\pi}(x)=[\pi_{ij}(x)]_{ij}$ is the predicted matrix of the posterior marginal probabilities of the edge occupancy for some model $M$. The calculation in Appendix E shows that, provided that the data is generated with $M^{*}$ and the reconstruction is performed with $M$, the posterior entropy and the expected posterior loss are equal if $M$ is equal to $M^{*}$ in distribution. Consequently, the reconstructability is linearly related to the posterior loss as follows:
@@
Fig. 4: Comparison between reconstructability and different performance metrics: (a) posterior loss (Eq. (20)), (b) mean error $\binom{N}{2}^{-1}\sum_{i<j}|a_{ij}-\pi_{ij}(x)|$, (c) area under the receiver operating characteristic curve (AUC) and (d) Jaccard similarity (see Ref. (Peixoto, 2019, Eq. 11)). Each point shows a different realization of the Glauber dynamics whose graphs are generated from the Erdős-Rényi model with $N=100$ nodes and $E=250$ edges, and whose initial conditions are random. Reconstructions are performed with the same model, whose parameters are fixed to those used for generating the data. We used time series of $T=500$ time steps (as in Fig. (2), the parameters other than the coupling constant $J$ are specified in Table 2). We generated 24 realizations of the process for each value of $J$ and used 30 different coupling values uniformly spaced between 0 and 0.5. These coupling values are fixed during inference. The colors indicated in the legend show the value of $J$ associated with the point (only 6 colors are shown for conciseness). Finally, we show the determination coefficients $R^{2}$ relating the performance metrics to $U(G^{*}|X^{*})$ in each plot. For panel (a), we used Eq. (21) directly to evaluate the determination coefficient, and for panels (b) and (d), we used standard linear regression to find the slope and estimate $R^{2}$. For panel (c), because the scaling is not linear like the other cases, we used instead log-linear regression to estimate $R^{2}$.
@@
III.4 Reconstructability of hierarchical Bayesian models
@@
We first consider the case where a TDG model includes hyperparameters, denoted $\theta^{*}$. Let $\theta$ and $G$ be the reconstructed parameters and graph, respectively, which are reconstructed via the posterior $P(\theta,G;X)$. The distribution $P(\theta^{*},G^{*})$ is generally different than the reconstruction model prior, i.e., $P(\theta,G)$. We consider the mutual information between the true variables $(\theta^{*},G^{*})$ and the reconstructed ones $(\theta,G)$ as the performance measure of interest, given again that $X$ is fixed by realizations of $X^{*}$. Since $(\theta^{*},G^{*})$ and $(\theta,G)$ are related through the condition $X=X^{*}$, the mutual information between them is bounded using the following data processing inequality:
$I(\theta^{*},G^{*};\theta,G|X=X^{*})\leq I(\theta^{*},G^{*};X^{*})\,.$ (23)
@@
IV Data-driven reconstructability and model selection
Until now, we have assumed that the TDG model $M^{*}$ was known. This allowed the mutual information $I(G^{*};X^{*})$, the upper bound of the performance metric $I(G^{*};G|X=X^{*})$, to be within reach. Outside of theoretical settings, the TDG process is generally unknown. Hence, we generally cannot evaluate the true reconstruction limit, although we may have access to many realizations of $X^{*}$ which should help get closer to it. Three remarks are in order. First, reconstructability is independent of the observations; it strictly depends on $M^{*}$. Second, all reconstruction models $M$ have a reconstructability, noted $U(G|X)$, which is calculated identically to Eq. (17). However and thirdly, the reconstructability of $M$ differ in two ways from $U(G^{*}|X^{*})$: (i) their values are potentially different and (ii) the data generation process is $M^{*}$, not $M$. Consequently, it may be possible to leverage the reconstructability of $M$, with these considerations in mind, to get a data-driven proxy of the true upper bound $U(G^{*}|X^{*})$.
IV.1 Reconstruction index based on information gain
To bring back the dependency of the reconstructability on the observations, we take a similar approach as before and start with an information measure. For a model $M$, the data-driven version of mutual information is called the information gain Mitchell, T. M. (1997), and it is defined as
@@ IV.1 Reconstruction index based on information gain
$=\sum_{g\in\mathcal{G}}P(G=g|X=x)\log\left(\frac{P(G=g|X=x)}{P(G=g)}\right)\,,$ (25)
which may depend on any realization $x\in\mathcal{X}$. Note that the expectation of the information gain yields back the mutual information between $G$ and $X$, i.e., $\mathbb{E}_{X}\quantity[\mathcal{I}_{M}(X)]=I(G;X)$. The information gain measures the reduction in the entropy of a variable $G$ achieved by learning the state $x$ of another variable $X$. It is primarily used in feature selection, especially decision tree training, where it is used as a criterion for how to best split the data (Mitchell, T. M., 1997, Chapter 3). Like the mutual information, the information gain can be shown to be non-negative (see Appendix F) and upper-bounded:
$0\leq\mathcal{I}_{M}\leq\Lambda_{M}$
@@
IV.2 Interpretation of the reconstruction index under incorrect assumptions
@@
We must be careful in our interpretation of the reconstruction index, as its value can be misleading if not used in the correct way. Figure 5 shows how the reconstruction index behaves when the reconstruction model is incorrect to different extents. In this example, we generated time series of the Glauber dynamics with a given coupling constant $J^{*}$, and reconstructed the graphs using the same Glauber model, but typically with an erroneous coupling constant $J\neq J^{*}$. As we can see in Fig. 5(a), the reconstruction index keeps increasing as $J$ gets larger, even when it gets larger than $J^{*}$. While the reconstruction index is larger for $J>J^{*}$, the posterior loss actually shows, as expected, that the reconstruction becomes worse. Figure 5(a) also illustrates how the reconstruction index correlates with the posterior loss, depending on $J$. Indeed, as $J$ gets closer to $J^{*}$, the reconstruction index converges to the true reconstructability of the reconstruction model, which increasingly becomes linearly related to the posterior loss as previously shown (see Eq. (21) and Appendix E). In this regime, the reconstruction index is a good proxy of the true reconstructability because $M$ properly approximates the behavior of $M^{*}$.
The behavior of the reconstruction index when the reconstruction model is incorrect raises some important remarks. Recall that, fundamentally, the reconstruction index is a normalized version of the KL divergence between the posterior and the prior of the reconstruction model. Therefore, it is perhaps not surprising that we lose the correspondence between reconstruction index and performance established in Sec. III.3 when the model is incorrect. Indeed, having a posterior that is very different from the prior implies a high reconstruction index even though the posterior distribution is actually wrong.
@@
IV.3 Role of evidence-based model selection

### v2

@@ II.2.1 Entropy
II.2.2 Network reconstructability
Those information-theoretic tools can be used to define the reconstruction limit. Consider the mutual information between the true and the reconstructed random graphs
$I(G^{*};\hat{G})=H(G^{*})-H(G^{*}|\hat{G})\,,$ (11)
where
$H(G^{*}|\hat{G})=-\mathbb{E}_{G^{*},\hat{G}}\quantity[\log P(G^{*}|\hat{G})]\,$ (12)
is the entropy of the true graph given the reconstructed one. The conditional probability $P(G^{*}|\hat{G})=P(G^{*},\hat{G})/P(\hat{G})$ is such that both $P(G^{*},\hat{G})$ and $P(\hat{G})$ are marginal distributions of $P(G^{*},X^{*},\hat{G})$ [Eq. (2)]. Three observations regarding this performance measure are in order. First, the quantity $I(G^{*};\hat{G})$ may be interpreted as measuring the similarity between the information contents of $G^{*}$ and $\hat{G}$. The higher it is, the more similar $G^{*}$ and $\hat{G}$ are and the better is the reconstruction. Conversely, when $I(G^{*};\hat{G})=0$, it is minimized and both graphs are independent from one another. Note that similar mutual information measures have been used as a performance measure in the context of community detection for comparing pairs of partitions Newman et al. (2020); Jerdee et al. (2024).
Second, $I(G^{*};\hat{G})$ is related to the probability of error, defined as
$p_{e}=P(\epsilon)\,,$ (13)
where $\epsilon=\mathbb{I}[G^{*}\neq\hat{G}]$, with $\mathbb{I}[\cdots]$ being the indicator function, denotes a binary random variable that takes the value 1 when $G^{*}\neq\hat{G}$ and 0 otherwise. This relationship can be shown through Fano’s inequality Cover and Thomas (2006):
$H(G^{*}|\hat{G})\leq h(p_{e})+H(G^{*})p_{e}\,,$ (14)
where $h(p)\equiv-p\log p-(1-p)\log(1-p)$ is the binary entropy. Indeed, given that $h(p_{e})\leq 1$, modifying Fano’s inequality yields
$p_{e}\geq 1-\frac{I(G^{*};\hat{G})+1}{H(G^{*})}\,.$ (15)
This lower bound on the probability of error is minimized when $I(G^{*};\hat{G})$ is maximized.
Third, using the data processing inequality Cover and Thomas (2006), it is also related to the mutual information between $G^{*}$ and $X^{*}$ as follows:
$I(G^{*};\hat{G})\leq I(G^{*};X^{*})\,,$ (16)
where the mutual information upper bound is expressed as
@@ II.2.2 Network reconstructability
is the mutual information between the true graph $G^{*}$ and the data process $X^{*}$. Intuitively, $I(G^{*};X^{*})$ quantifies the amount of reconstructible information that both $X^{*}$ and $G^{*}$ share—i.e., the amount of information that $X^{*}$ contains about $G^{*}$ [see Fig. 1(b)]. The mutual information $I(G^{*};X^{*})$ also sets the maximum in reconstruction performance as measured by $I(G^{*};\hat{G})$: It is the reconstruction limit.
The mutual information $I(G^{*};X^{*})$ is itself bounded between 00 and $H(G^{*})$ Cover and Thomas (2006). When $I(G^{*};X^{*})=0$, $X^{*}$ and $G^{*}$ are independent and thus the data $X^{*}$ contains no information about the graph $G^{*}$. In turn, it is impossible for any reconstruction model $M$ to extract information from the data, regardless of its specification. When $I(G^{*};X^{*})=H(G^{*})$, the data $X^{*}$ contains all the information about the graph $G^{*}$. In this case, it is in principle possible to perfectly reconstruct the graph without any error, assuming the model $M$ is optimal.
@@
The reconstructability has been described thoroughly in Ref. Murphy et al. (2024) and helped unveiling a special duality between our ability to predict the time evolution of a system and our ability to reconstruct the interactions between its constituents. As it is a normalized version of the mutual information upper bound $I(G^{*};X^{*})$, the reconstructability is bounded between 00 and $1$. When $U(G^{*}|X^{*})=0$, any attempt at reconstruction is futile, whereas it is theoretically possible to decode all the information when $U(G^{*}|X^{*})=1$. As such, the reconstructability is a measure of the average proportion of information that can be extracted from the data about the graph. For instance, when it is equal to $\frac{1}{2}$, it precisely means that half of the graph information is, on average, contained in the data and that in turn half of it can possibly be reconstructed. We stress that $U(G^{*}|X^{*})=\frac{1}{2}$ may not be directly interpreted as half of the graph’s edges being reconstructible. Rather, information may generally be distributed in a heterogeneous way over the graph’s structure, as a single bit of information may reconstruct more than one edge in the graph depending on how correlated they are.
Going back to the earlier example of a single edge of Sec. II.1.2, we can perform the complete calculation analytically. First, we can calculate the entropy of the prior,
@@
This leads to the reconstructability of the edge using Eq. (18), which is plotted in Fig. 3. As expected, the reconstructability typically increases as the number of observations $T$ increases, even reaching 1 in some cases, e.g., when $q\to 1$. Also, notice how the reconstructability is zero for every value of $T$ when the true positive probability $q$ is equal to false positive probability $r$. This shows that as true positives and false positives become indistinguishable, the edge becomes impossible to reconstruct.
II.2.3 Optimal reconstruction performance
The reconstruction limit corresponds to the maximum performance, as measured by $I(G^{*};\hat{G})$, achievable by any algorithm. Hence, any reconstruction model that is capable of reaching this limit, i.e., $I(G^{*};\hat{G})=I(G^{*};X^{*})$, is said optimal in its reconstruction abilities. It is not surprising that the TDG model is optimal according to this definition. As a result, the reconstructability $U(G^{*}|X^{*})$ can also be interpreted as a reconstruction performance measure of the TDG model $M^{*}$.
In fact, the reconstructability relates to standard performance measures. One such example is the posterior loss—also known as the log loss and the cross-entropy loss in the machine learning community. This measure is defined as
@@ II.2.3 Optimal reconstruction performance
where $\bm{a}^{*}$ denotes the adjacency of the true graph $g^{*}$, such that $a_{ij}^{*}$ counts the number of edges connecting the nodes $i$ and $j$ (we use the convention that $a_{ii}^{*}$ is always a multiple of 2) and $\bm{\pi}(x)=[\pi_{ij}(x)]_{ij}$ is the predicted matrix of the posterior marginal probabilities of the edge occupancy for some model $M$. We show in Appendix IV.4 that, provided that the data is generated with $M^{*}$ and the reconstruction is performed with $M$, the posterior entropy and the expected posterior loss are equal if $M$ is equal to $M^{*}$ in distribution. Consequently, the reconstructability is linearly related to the posterior loss as follows:
@@
Fig. 4: Comparison between reconstructability and different performance metrics: (a) posterior loss (Eq. (21)), (b) mean error $\binom{N}{2}^{-1}\sum_{i<j}|a_{ij}-\pi_{ij}(x)|$, (c) area under the receiver operating characteristic curve (AUC) and (d) Jaccard similarity (see Ref. (Peixoto, 2019, Eq. 11)). Each point shows a different realization of the Glauber dynamics whose graphs are generated from the Erdős-Rényi model with $N=100$ nodes and $E=250$ edges, and whose initial conditions are random. Reconstructions are performed with the same model, whose parameters are fixed to those used for generating the data. We used time series of $T=500$ time steps (as in Fig. 2, the parameters other than the coupling constant $J$ are specified in Table 2). We generated 24 realizations of the process for each value of $J$ and used 30 different coupling values uniformly spaced between 0 and 0.5. These coupling values are fixed during inference. The colors indicated in the legend show the value of $J$ associated with the point (only 6 colors are shown for conciseness). Finally, we show the determination coefficients $R^{2}$ relating the performance metrics to $U(G^{*}|X^{*})$ in each plot. For panel (a), we used Eq. (22) directly to evaluate the determination coefficient, and for panels (b) and (d), we used standard linear regression to find the slope and estimate $R^{2}$. For panel (c), because the scaling is not linear like the other cases, we used instead log-linear regression to estimate $R^{2}$.
@@
II.2.4 Reconstructability of hierarchical Bayesian models
@@
Consider the case where a TDG model with variables $(G^{*},X^{*})$ also includes hyperparameters, denoted $\theta^{*}$ with probability distribution $P(\theta^{*})$, such that $G^{*}$ is conditioned on $\theta^{*}$—i.e., $P(\theta^{*},G^{*})=P(\theta^{*})P(G^{*}|\theta^{*})$. Let $\hat{\theta}$ and $\hat{G}$ be the reconstructed random parameters and graph, respectively, which are reconstructed from $X^{*}$ via some distribution $P(\hat{\theta},\hat{G}|X^{*})$. The random variables $(\theta^{*},G^{*})$ are related to those of the reconstruction model $(\hat{\theta},\hat{G})$ via $X^{*}$ as follows:
$P(\theta^{*},G^{*},X^{*},\hat{\theta},\hat{G})=P(\theta^{*},G^{*})P(X^{*}|G^{*})P(\hat{\theta},\hat{G}|X^{*})\,,$ (24)
where, again assuming that we use a Bayesian reconstruction model $M$, we let $P(\hat{\theta}=\vartheta,\hat{G}=g|X^{*}=x^{*})=P(\theta=\vartheta,G=g|X=x^{*})$, which is given by Eq. (23). In this case, the mutual information between $(\theta^{*},G^{*})$ and $(\hat{\theta},\hat{G})$ can be bounded using the following data processing inequality:
$I(\theta^{*},G^{*};\hat{\theta},\hat{G})\leq I(\theta^{*},G^{*};X^{*})\,.$ (25)
@@
II.3 Data-driven reconstructability and model selection
Until now, we have assumed that the TDG model $M^{*}$ was known to compute the mutual information $I(G^{*};X^{*})$. Outside of theoretical settings however, the TDG process is typically unknown. Hence, we generally cannot evaluate the true reconstruction limit, although we may have access to many realizations of $X^{*}$ which should help get closer to it. Three remarks are in order. First, the reconstructability is independent of the observations; it strictly depends on $M^{*}$. Second, any generative model $M$—unlike nongenerative ones—has a reconstructability, noted $U(G|X)$, which is calculated identically to Eq. (18). In other words, the condition that $M$ is capable of generating new data is crucial to our ability to calculate a reconstructability value. However and thirdly, the reconstructability $U(G|X)$ of $M$ differ in two ways from $U(G^{*}|X^{*})$ related to the actual reconstruction limit of the data: (i) their values are potentially different and (ii) the data generation process is $M^{*}$, not $M$. Consequently, we can leverage the reconstructability of $M$, with these considerations in mind, to get a data-driven proxy of the true upper bound $U(G^{*}|X^{*})$.
II.3.1 Reconstruction index based on information gain
To bring back the dependency of the reconstructability on the observations, we take a similar approach as before and start with an information measure. For a model $M$ and any instance $x\in\mathcal{X}$, the data-driven version of mutual information is called the information gain Mitchell, T. M. (1997), and it is defined as
@@ II.3.1 Reconstruction index based on information gain
$=\sum_{g\in\mathcal{G}}P(G=g|X=x)\log\left(\frac{P(G=g|X=x)}{P(G=g)}\right)\,.$ (27)
Note that the expectation of the information gain yields back the mutual information between $G$ and $X$, i.e., $\mathbb{E}_{X}\quantity[\mathcal{I}_{M}(X)]=I(G;X)$. The information gain measures the reduction in the entropy of a variable $G$ achieved by learning the state $x$ of another variable $X$. It is primarily used in feature selection, especially decision tree training, where it is used as a criterion for how to best split the data (Mitchell, T. M., 1997, Chapter 3). Like the mutual information, the information gain can be shown to be non-negative (see Appendix IV.5) and upper-bounded:
$0\leq\mathcal{I}_{M}\leq\Lambda_{M}\,,$
@@
II.3.2 Interpretation of the reconstruction index under incorrect assumptions
@@
We must be careful in our interpretation of the reconstruction index, as its value can be misleading if not used in the correct way. Figure 5 shows how the reconstruction index behaves when the reconstruction model is incorrect to different extents. In this example, we generated time series of the Glauber dynamics with a given coupling constant $J^{*}$, and reconstructed the graphs using the same Glauber model, but typically with an erroneous coupling constant $J\neq J^{*}$. As we can see in Fig. 5(a), the reconstruction index keeps increasing as $J$ gets larger, even when it gets larger than $J^{*}$. While the reconstruction index is larger for $J>J^{*}$, the posterior loss actually shows, as expected, that the reconstruction becomes worse. Figure 5(a) also illustrates how the reconstruction index correlates with the posterior loss, depending on $J$. Indeed, as $J$ gets closer to $J^{*}$, the reconstruction index converges to the true reconstructability of the reconstruction model, which increasingly becomes linearly related to the posterior loss as previously shown (see the Appendix IV.4) In this regime, the reconstruction index is a good proxy of the true reconstructability because $M$ properly approximates the behavior of $M^{*}$.
The behavior of the reconstruction index when the reconstruction model is incorrect raises some important remarks. Recall that, fundamentally, the reconstruction index is a normalized version of the KL divergence between the posterior and the prior of the reconstruction model. Therefore, it is perhaps not surprising that we lose the correspondence between reconstruction index and performance established in Sec. II.2.3 when the model is incorrect. Indeed, having a posterior that is very different from the prior implies a high reconstruction index even though the posterior distribution is actually wrong.
@@
II.3.3 Role of evidence-based model selection